OpenStudy (anonymous):

In a simple pendulum graph, the value of the slope is equivalent to : 1.g(acceleration due to gravity) 2.L(length) 3.T(period) 4.Other(specify)

OpenStudy (astrophysics):

The pendulum graphs are often period squared vs. length, so if you put it together you should definitely get gravity.

OpenStudy (astrophysics):

Note that the slope is \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} } = \frac{ \Delta y }{ \Delta x }\]

OpenStudy (michele_laino):

since, the formula for period T of a simple pendulum is: \[\Large T = 2\pi \sqrt {\frac{L}{g}} \] we can write: \[\Large {T^2} = \frac{{4{\pi ^2}}}{g}L\] so, if we make a plot where along the vertical axis, there are the values of T^2, whereas along the horizontal axis there are the values of the length L of the pendulum, then such plot is a straight line whose slope is the subsequent quantity: \[\Large \frac{{4{\pi ^2}}}{g}\]